Ch31 P7 What is the radius of 238U? Enter with 3SigFigs and proper units.
Answer #1

Ans-

We have radius of nucleus formula,

R= Ro(A)^(1/3), Here A= Mass number = 238

Ro= 1.2*10^(-15)m

So, R= 1.2*10^(-15) * ((238)^(1/3))m= 7.43*10^(-15) m

P.S : Please do give a thumbs up since we as an expert take pain and take our time out by typing answer in lucid language and putting best of our knowledge at your doorstep.
Thank you